OpenStudy (df001):

system of equation help 2x+y-3z=-12 3x-2y-z=3 -x+5y+2z=-3

OpenStudy (df001):

Please help me!!! I started with the third equation and got x=5y+2z+3

OpenStudy (df001):

then i substitued it to the first and second equation and got 11y-1z=-6 and 13y+5z=-6

OpenStudy (baru):

continue the same way, with your two new equations

OpenStudy (baru):

from the first equation: z=11y + 6 substitute for z in 13y+5z=-6

OpenStudy (df001):

is z=4??

OpenStudy (anonymous):

Yes

OpenStudy (df001):

Z being 4, I replaced it in the equation 13y+5z=-6 and got y=-2

OpenStudy (df001):

then 5y+2z+3=x i replaced y=-2 and z=4 and got x=-9

OpenStudy (df001):

Please confirm my answers y=-2 z=4 x=-9